OpenStudy (anonymous):

I need help with Interval notation of X >-2

OpenStudy (anonymous):

\[(-2,\infty)\]

OpenStudy (saifoo.khan):

\[\Huge \color\red {-2, \infty}\]

OpenStudy (anonymous):

Make sure to read your text book! :))

OpenStudy (anonymous):

forgot something didn't you?

OpenStudy (anonymous):

\[\color{#ff0099}{\huge(-2,\infty)}\]

OpenStudy (anonymous):

this is getting silly

OpenStudy (saifoo.khan):

\begin{array}l\color{red}{\text{{}}\color{orange}{\text{−}}\color{#9c9a2e}{\text{2}}\color{green}{\text{,}}\color{blue}{\text{∞}}\color{purple}{\text{}}}\color{purple}{\text{}}\end{array}

OpenStudy (anonymous):

oops

OpenStudy (saifoo.khan):

\[\Huge \begin{array}l\color{red}{\text{−}}\color{orange}{\text{2}}\color{#9c9a2e}{\text{,}}\color{green}{\text{∞}}\color{blue}{\text{}}\end{array}\]

OpenStudy (anonymous):

lol \[L_AT^EX\]